OpenStudy (ttp):

Which expression is equivalent to 5^-4 over 5 ^1 5^-6 5^-5 5^-4 5^-3

OpenStudy (ranga):

\[\frac{ x ^{m} }{ x ^{n} } = x ^{m - n}\] Will that help you solve the problem?

OpenStudy (anonymous):

ah should have tagged me haha

OpenStudy (anonymous):

\[\frac{ 5^{-4} }{ 5 }\]

OpenStudy (anonymous):

So that's your question?

OpenStudy (ttp):

that's the question:)

OpenStudy (anonymous):

omg lol k sec

OpenStudy (ttp):

lol

OpenStudy (anonymous):

\[\frac{ 1 }{ 5^{4}*5^1 }\]

OpenStudy (anonymous):

so...

OpenStudy (anonymous):

\[\frac{ 1 }{ 5^5 }\]

OpenStudy (ttp):

B?

OpenStudy (anonymous):

\[5^{-5}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Was showing it step by step but yes b :)

OpenStudy (ttp):

lol :)